OpenStudy (anonymous):

Determine the angular displacement in radians for 4.3 revolutions

OpenStudy (anonymous):

1 revolution = 2pi in radians you have 4.3 revolutions. What do you do?

OpenStudy (anonymous):

4.3 / 2?

OpenStudy (anonymous):

4.3/2pi**** don't forget the pi. very crucial.

OpenStudy (anonymous):

6.7?

OpenStudy (anonymous):

how did you get 6.7

OpenStudy (anonymous):

your equation should look as so:\[\frac{ 4.3 }{2\pi}\]

OpenStudy (anonymous):

.7?

OpenStudy (anonymous):

Yes. If you round it up. Good job.

OpenStudy (selrachcw95):

Have u tried Kahn academy?

OpenStudy (selrachcw95):

Just wondering.

OpenStudy (anonymous):

no i haven't

OpenStudy (haseeb96):

Formula of Angular diaplacement is \[\theta=\frac{ s }{ r }\] s is the distance covered by the object R is radius SO circle is 2phi 4.3\2 x 3.14 since phi=3.141 4.3\6.28 0.6847 0.7 approximately this is the correct answer